OpenStudy (surryyy):

Solve the equations by combining like terms. Show all the steps 16x − 4x = -48

OpenStudy (exoexoexo):

16x-4x=12x 12x=-48 12x/12=-48/12 x=-4

OpenStudy (surryyy):

ty

OpenStudy (exoexoexo):

np